# Spatial-Aware Conformal Prediction for Trustworthy Hyperspectral Image Classification

### v1

@@ Abstract
Hyperspectral image (HSI) classification involves assigning specific labels to each pixel to identify various land cover categories. Although deep classifiers have shown high predictive accuracy in this field, quantifying their uncertainty remains a significant challenge, which hinders their application in critical contexts. This study first theoretically evaluates the applicability of Conformal Prediction (CP), an emerging technique for uncertainty quantification, in the context of HSI classification. We then propose a conformal procedure that provides HSI classifiers with trustworthy prediction sets, offering coverage guarantees that ensure these sets contain the true labels with a user-specified probability. Building on this foundation, we introduce Spatial-Aware Conformal Prediction (SACP), which incorporates essential spatial information inherent in HSIs by aggregating non-conformity scores of pixels with high spatial correlation. Both theoretical and empirical results demonstrate that SACP outperforms standard CP in HSI classification. The source code is accessible at https://github.com/J4ckLiu/SACP.
@@ I Introduction
Hyperspectral imaging technology captures spectral information across a broad wavelength range, providing more detailed pixel features than traditional color images [1, 2], which are limited to three channels. Hyperspectral image (HSI) classification is a fundamental task involving the assignment of discrete labels to each pixel to delineate various land cover categories. Despite the high accuracy achieved by deep classifiers in HSI classification [3, 4, 5, 6, 7], the quantification of their uncertainty remains inadequately explored and presents a substantial challenge. Effective uncertainty quantification is imperative due to the diverse and significant practical applications of HSIs, including surveillance, threat detection [8], mineral detection [9], agriculture [10], environmental protection [11], and defense [12], where errors could lead to serious consequences. For instance, in threat detection or mineral exploration, incorrect classifications could result in misidentifying potential threats or valuable resources, which can have severe repercussions. These applications highlight the urgent need to quantify uncertainty in deep HSI classifier decisions and to advance related methodologies.
Various techniques have been developed to estimate and incorporate uncertainty into predictive models, including confidence calibration [13], Monte-Carlo Dropout [14], and Bayesian neural networks [15]. Although these techniques provide valuable insights into decision uncertainty, they lack theoretical guarantees regarding their effectiveness. This highlights the significance of Conformal Prediction (CP) [16, 17, 18]. CP is a statistical framework that can be applied to any predictive model. Instead of predicting a single label based on the output probabilities of a classifier, this framework generates prediction sets that guarantee the inclusion of the true label with a user-specified coverage probability. Thus, it provides a formal performance guarantee, as follows:
$\mathbb{P}(Y_{\text{test}}\in\mathcal{C}(X_{\text{test}})))\geq 1-\alpha,$ (1)
where $Y_{\text{test}}$ is the true label, $\mathcal{C}(X_{\text{test}})$ is the prediction set, and $\alpha$ is the user-specified error rate. The strong theoretical foundation of CP supports its efficacy across a variety of practical applications, ensuring both reliable and trustworthy decision-making. Conformal predictors may generate prediction sets of size 1 for highly confident predictions and larger sets for those with greater uncertainty. Subsequently, human experts can make the final judgments for the predictions sets with larger sizes. Consequently, CP is a highly suitable candidate for quantifying the uncertainty of deep HSI classifiers.
@@
The application of CP to HSI classification remains unexplored, with two primary considerations that need to be addressed for its effective implementation. Firstly, unlike traditional classification tasks, HSI classification models require access to the entire HSI during their training. This overlap could potentially undermine the validity of CP, which typically relies on the exchangeability of calibration and test data. It is therefore imperative to theoretically examine whether the validity of CP is maintained in the context of HSI classification. Secondly, the performance of CP algorithms is dictated by the non-conformity score function once the base classifier is determined. Although many effective non-conformity score functions have been proposed—such as Adaptive Prediction Sets (APS) [19], which generates prediction sets by accumulating descending softmax values, and Regularized Adaptive Prediction Sets (RAPS) [20], which enhances performance through regularization based on class ranks—these score functions generally neglect the critical spatial information inherent in HSIs. This oversight results in suboptimal prediction sets with unnecessarily large sizes. In summary, addressing the following two questions is essential for the effective application of CP in HSI classification:
  • Does the validity of CP persist in the context of HSI classification?
  • How can spatial information be leveraged to enhance the performance of CP in HSI classification?
@@
In this paper, we first undertake a theoretical analysis to confirm the validity of applying CP to HSI classification, even when classifiers are exposed to pixels from both the calibration and test sets during training. Building upon this foundation, we propose a framework for integrating conformal prediction sets into HSI classifiers. We then theoretically demonstrate that incorporating neighborhood information can enhance the statistical efficiency of prediction sets. Motivated by this, we introduce Spatial-Aware Conformal Prediction (SACP), which refines CP by aggregating non-conformity scores of pixels with high spatial correlation. This approach effectively leverages the spatial information inherent in HSIs to improve overall CP performance. Additionally, we provide theoretical demonstration that SACP guarantees coverage.
@@
We conduct extensive experiments on standard HSI classification benchmarks, including Indian Pines, Pavia University, and Salinas. The experimental results indicate that SACP significantly improves the efficiency of various non-conformity score functions in HSI classification, outperforming standard CP while maintaining satisfactory coverage rates across diverse scenarios. Furthermore, SACP demonstrates its effectiveness with both conventional classifiers, such as CNN-based models, and advanced classifiers, including Transformer-based models, highlighting the versatility of our proposed approach.
Our main contributions are summarized as follows:
  • To the best of our knowledge, we are the first to investigate the application of CP to HSI classification. We conduct a theoretical analysis to evaluate the validity of employing CP for this fundamental task.
@@
  • We introduce a framework for equipping HSI classifiers with reliable conformal prediction sets, which ensure the inclusion of true labels with a user-specified probability.
@@
  • We provide a theoretical demonstration that incorporating spatial information can enhance the statistical efficiency of prediction sets. Based on this insight, we propose SACP, a framework that leverages the spatial information inherent in HSIs to improve CP performance. We also theoretically establish that SACP ensures coverage.
@@
  • Extensive experiments validate the effectiveness of our proposed method. The results show that SACP outperforms standard CP in HSI classification by reducing the average size of prediction sets while maintaining satisfactory coverage rates across various scenarios.
@@ II-A Conformal Prediction
$\hat{\pi}_{\theta,y}(\bm{x})=\sigma(f_{\theta}(\bm{x}))_{y}=\frac{e^{f_{\theta,y}(\bm{x})}}{\sum_{i=1}^{\mathit{K}}e^{f_{\theta,i}(\bm{x})}},$ (2)
@@
Pipeline of Conformal Prediction
Instead of predicting a single label based on the output probabilities of a classifier, CP constructs a set-valued mapping $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ that satisfies the following marginal coverage property:
$\mathbb{P}(\mathit{Y}\in\mathcal{C}(\mathit{X}))\geq 1-\alpha,$ (3)
@@
Before deployment, CP starts with a calibration step using a separate calibration set, $\mathcal{D}_{\text{cal}}:=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, drawn independently from the distribution $\mathcal{P}_{\mathcal{XY}}$. Specifically, we compute the non-conformity score $s_{i}=\mathit{S}(\bm{x}_{i},y_{i})$ for each sample in $\mathcal{D}_{\text{cal}}$, which measures how $y_{i}$ “conforms” to the prediction at $\bm{x}_{i}$. The $1-\alpha$ quantile of the scores $\{s_{i}\}_{i=1}^{n}$ is then determined as the threshold $\tau$. Formally, $\tau$ is obtained as follows:
@@
where $\mathcal{C}_{1-\alpha}(\bm{x}_{n+1};\tau)$ includes all labels for which their non-conformity score $\mathit{S}(\bm{x}_{n+1},y)$ does not exceed the threshold $\tau$. This equation illustrates a nesting property of the threshold. Specifically, if $\tau_{1}\leq\tau_{2}$, then
$\mathcal{C}_{1-\alpha}(\bm{x}_{n+1};\tau_{1})\subseteq\mathcal{C}_{1-\alpha}(\bm{x}_{n+1};\tau_{2}).$ (6)
As stated in the following theorem, the prediction sets generated by Eqs. (4) and (5) are guaranteed to achieve finite-sample marginal coverage of the true labels.
@@
Once the base classifier is determined, the effectiveness of the resulting prediction set depends critically on the chosen non-conformity score function. Two popular options are Adaptive Prediction Sets (APS) and Regularized Adaptive Prediction Sets (RAPS). APS calculates the non-conformity scores by accumulating the softmax probabilities, sorted in descending order:
$S_{\text{APS}}(\bm{x},y)\!=\!\sum_{y_{i}\neq y}\hat{\pi}_{\theta,y_{i}}(\bm{x})\cdot\mathds{1}_{\{\hat{\pi}_{\theta,y_{i}}(\bm{x})>\hat{\pi}_{\theta,y}(\bm{x})\}}+u\cdot\hat{\pi}_{\theta,y}(\bm{x}),$ (7)
where $u$ is an independent random variable following a uniform distribution on $[0,1]$. However, the softmax probabilities typically exhibit a long-tailed distribution, which facilitates the inclusion of tail classes in the prediction sets. To address this issue, RAPS [20] refines prediction sets by incorporating regularization for the classes. Formally, the non-conformity score function of RAPS is defined as follows:
$\begin{array}[]{r}S_{\textrm{RAPS}}(\bm{x},y)=S_{\textrm{APS}}(\bm{x},y)+\lambda\cdot(o(y,\hat{\pi}_{\theta}(\bm{x}))-k)^{+},\end{array}$ (8)
where $o(y,\hat{\pi}_{\theta}(\bm{x}))$ is the label ranking of $y$, $\lambda$ and $k$ are hyperparameters, and $(z)^{+}$ denotes the positive part of $z$. SAPS [21] reduces the emphasis on probabilities by retaining only the maximum softmax probability and discarding the others. Formally, SAPS is defined as follows:
$S_{\text{SAPS}}(\bm{x},y)=\left\{\begin{array}[]{l}u\cdot\hat{\pi}_{\theta,\max}(\bm{x}),\quad\text{if }o(y,\hat{\pi}_{\theta}(\bm{x}))=1,\\ \hat{\pi}_{\theta,\max}(\bm{x})+(o(y,\hat{\pi}_{\theta}(\bm{x}))-2+u)\cdot\lambda,\text{else},\end{array}\right.$ (9)
where $\lambda$ balances the weight of ranking information, $\hat{\pi}_{\theta,\max}(\bm{x})$ denotes the maximum softmax probability of instance $\bm{x}$, and $u$ is a random variable uniformly distributed on $[0,1]$. It can be concluded that APS, RAPS, and SAPS depend exclusively on softmax probabilities and label ranks.
@@
$\text{Coverage}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{(\bm{x},y)\in\mathcal{D}_{\text{test}}}\mathds{1}_{\{y\in\mathcal{C}(\bm{x})\}},$ (10)
which should ideally match $1-\alpha$. Size measures the average length of prediction sets. When coverage is satisfied, a smaller size indicates higher statistical efficiency and is more desirable. Formally, Size is defined as:
$\text{Size}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{(\bm{x},y)\in\mathcal{D}_{\text{test}}}|\mathcal{C}(\bm{x})|.$ (11)
SSCV [20] effectively reflects the conditional coverage rate and is formally defined as:
$\text{SSCV}=100\times\sup_{j}\left|(1-\alpha)-\frac{\left|\{i:y_{i}\in\mathcal{C}(\bm{x}_{i})\text{ and }i\in\mathcal{J}_{j}\}\right|}{|\mathcal{J}_{j}|}\right|,$ (12)
@@
II-B Related Works
Despite the high accuracy achieved by deep classifiers in HSI classification, the quantification of their uncertainty remains inadequately explored and presents a substantial challenge. Uncertainty quantification techniques applied to graph-based HSI classifiers have been investigated in [22]. This study adapts advanced uncertainty quantification models, originally designed for node classification in graphs, to the context of HSI classification. Additionally, prior work [23] utilizes Bayesian deep learning to estimate prediction uncertainty while also enhancing the structural characteristics of the models. However, these techniques lack theoretical guarantees, which underscores the importance of CP [16]. Typically, CP is applied in an inductive manner, also known as split conformal prediction, which employs a separate calibration set to improve computational efficiency. Notable variations include cross-validation methods [16] and the jackknife technique [24]. CP has demonstrated its versatility across various fields, including classification [25, 21], regression [26], and specialized areas such as image generative models [27], robotic control [28], graph neural networks [29, 30], and large language models (LLMs) [31, 32, 33]. Despite its broad application, the use of CP in HSIs remains unexplored, which motivates our research.
III Validity of CP on HSI
HSI Classification
Let $\mathcal{I}\in\mathbb{R}^{H\times W\times U}$ be the HSI to classify, where $H$ and $W$ denote the height and width in pixels, and $U$ denotes the number of spectral bands. Let $M=HW$ be the total number of pixels. Typically, HSI classification aims to predict the label $y\in\mathcal{Y}$ for each pixel in the image, given that a random subset of labels known. The pixels with unknown labels appear randomly at any location. When employing a deep classifier, the $(i,j)$-th pixel is generally considered as the center, with information from the small cuboid $B\in\mathbb{R}^{P\times P\times U}$ utilized to extract the joint spectral-spatial feature representation for the central pixel [34] [^1] In prior research, the terms “cuboids”, “cubes”, and “patches” are often used interchangeably to describe local regions of HSIs. For clarity, this work consistently uses the term “cuboids”. . Here, $P$ denotes the patch size of $B$. Correspondingly, the dataset constructed from $\mathcal{I}$ consists of $M$ such cuboids, as follows:
$\mathcal{D}_{\mathcal{I}}:=\{(B_{i},y_{i})\mid 1\leq i\leq M\},$ (13)

### v2

@@ Abstract
Hyperspectral image (HSI) classification involves assigning unique labels to each pixel to identify various land cover categories. While deep classifiers have achieved high predictive accuracy in this field, they lack the ability to rigorously quantify confidence in their predictions. Quantifying the certainty of model predictions is crucial for the safe usage of predictive models, and this limitation restricts their application in critical contexts where the cost of prediction errors is significant. To support the safe deployment of HSI classifiers, we first provide a theoretical proof establishing the validity of the emerging uncertainty quantification technique, conformal prediction, in the context of HSI classification. We then propose a conformal procedure that equips any trained HSI classifier with trustworthy prediction sets, ensuring that these sets include the true labels with a user-specified probability (e.g., 95%). Building on this foundation, we introduce Spatial-Aware Conformal Prediction (SACP), a conformal prediction framework specifically designed for HSI data. This method integrates essential spatial information inherent in HSIs by aggregating the non-conformity scores of pixels with high spatial correlation, which effectively enhances the efficiency of prediction sets. Both theoretical and empirical results validate the effectiveness of our proposed approach. The source code is available at https://github.com/J4ckLiu/SACP.
@@ I Introduction
Hyperspectral imaging technology captures spectral information across a wide wavelength range, providing more detailed pixel features compared to traditional three-channel color images [1, 2]. Hyperspectral image (HSI) classification, a fundamental task in this domain, involves assigning unique labels to each pixel to identify various land cover categories. While deep classifiers have demonstrated high accuracy in this field [3, 4, 5, 6, 7], performance metrics like accuracy alone are insufficient for their deployment in safety-critical contexts. In these scenarios, it is also crucial to quantify the certainty of predictions. This need is amplified by the significant applications of HSIs, including surveillance, threat detection [8], mineral detection [9], agriculture [10], environmental protection [11], and defense [12], where prediction errors could have serious consequences. For example, in threat detection or mineral exploration, incorrect predictions could result in misidentifying potential threats or valuable resources, leading to severe repercussions. This emphasizes the necessity of advancing methodologies that rigorously quantify the uncertainty of HSI classifiers to ensure their safe deployment.
Various techniques have been developed to estimate and incorporate uncertainty into predictive models, including confidence calibration [13], Monte Carlo Dropout [14], and Bayesian neural networks [15]. Additionally, there are uncertainty quantification strategies specifically tailored to HSI classification. For instance, PL-CNN [16] reduces both predictive and label uncertainty for more reliable predictions, achieving high classification accuracy while lowering the Expected Calibration Error (ECE) [13]. Another approach leverages Bayesian deep learning to estimate prediction uncertainty while enhancing the structural characteristics of HSI classifiers [17]. Furthermore, uncertainty quantification techniques specifically applied to graph-based HSI classifiers have been explored in [18]. This study adapts advanced uncertainty quantification models, initially developed for node classification in graphs, to the context of HSI classification. Collectively, these techniques contribute valuable insights into quantifying or reducing decision uncertainty. However, they lack theoretical guarantees of model performance [19], underscoring the importance of conformal prediction [20, 21, 22].
Unlike the techniques discussed earlier, conformal prediction—an emerging uncertainty quantification methodology—typically operates as post-hoc processing for the outputs of trained classifiers. This statistical framework can be applied to any predictive model, transforming the output probabilities into prediction sets while ensuring that the true label is included with a user-specified coverage probability. Notably, the validity of the prediction sets is assured in a distribution-free manner, providing explicit, non-asymptotic guarantees without reliance on any distributional or model assumptions [20]. Formally, this guarantee can be expressed as follows:
$\mathbb{P}(Y_{\text{test}}\in\mathcal{C}(X_{\text{test}}))\geq 1-\alpha,$ (1)
where $Y_{\text{test}}$ is the true label, $\mathcal{C}(X_{\text{test}})$ is the generated prediction set, and $\alpha$ is the user-specified error rate. The prediction set will consistently include the true labels of the inputs with the specified error frequency. The strong theoretical foundation of conformal prediction supports its efficacy across various applications, including classification [23, 19], regression [24], and specialized areas such as image generative models [25], robotic control [26], graph neural networks [27, 28], and large language models (LLMs) [29, 30, 31].
@@
Despite its success in numerous domains, conformal prediction has remained unexplored in the context of HSI data. There are two primary challenges that need to be addressed for the valid and effective implementation of conformal prediction in HSI classification. Firstly, unlike traditional classification tasks, HSI classifiers typically require access to the entire HSI, including test instances, during training. This practice could potentially undermine the validity of conformal prediction, which relies on the exchangeability of calibration and test data [20]. Therefore, it is imperative to theoretically examine whether the validity of conformal prediction is maintained in the context of HSI classification. Secondly, the performance of conformal prediction algorithms is dictated by the non-conformity score function once the base classifier is determined. Although many effective non-conformity score functions have been proposed—such as Adaptive Prediction Sets (APS) [32], which calculates the non-conformity scores by accumulating descending softmax values, and Regularized Adaptive Prediction Sets (RAPS) [33], which enhances the efficiency of conformal prediction sets through regularization based on label ranks—these non-conformity score functions generally neglect the critical spatial information inherent in HSIs. This oversight can lead to suboptimal prediction sets with unnecessarily large sizes. In summary, addressing the following two questions is essential for the effective application of conformal prediction in HSI classification:
  • Does the validity of conformal prediction persist in the context of HSI classification?
  • How can spatial information be utilized to enhance the performance of conformal prediction?
@@
In this paper, we first present a theoretical analysis that confirms the validity of conformal prediction in HSI classification, even when classifiers are exposed to pixels from both calibration and test sets during training. Building on this foundation, we propose a framework for integrating conformal prediction sets into HSI classifiers, ensuring the inclusion of the true label with a user-specified coverage probability. Furthermore, we theoretically illustrate that incorporating neighborhood information can enhance the statistical efficiency of prediction sets. Motivated by this insight, we introduce Spatial-Aware Conformal Prediction (SACP), which refines standard conformal prediction by aggregating the non-conformity scores of pixels exhibiting high spatial correlation. This approach effectively leverages the spatial information inherent in HSIs to improve the performance of conformal prediction. Moreover, we present a theoretical demonstration of the validity of SACP, confirming that it guarantees coverage as defined in Eq. (1).
@@
We conduct extensive evaluations on standard HSI classification benchmarks, including Indian Pines, Pavia University, and Salinas. The experimental results confirm the validity of our proposed conformal framework (standard conformal prediction) and SACP for HSI classification: the generated prediction sets consistently include the true label of data with a user-specified coverage probability. Furthermore, SACP significantly enhances the efficiency of various non-conformity score functions while maintaining satisfactory coverage rates across different scenarios, outperforming standard conformal prediction in this context. This finding underscores the critical importance of incorporating spatial information into conformal prediction. Notably, SACP demonstrates its effectiveness with both simple classifiers, such as CNN-based models, and advanced classifiers, including Transformer-based models, thereby highlighting the adaptability of our proposed approach.
Our main contributions are highlighted as follows:
  • To the best of our knowledge, we are the first to explore the application of the emerging uncertainty quantification technique, conformal prediction, to HSI data. We present a thorough theoretical proof that establishes the validity of conformal prediction for HSI classification, thereby opening a new avenue for the safe utilization of HSI.
@@
  • We introduce a framework that equips HSI classifiers with reliable conformal prediction sets, ensuring the inclusion of true labels with a user-specified probability.
@@
  • We provide a theoretical demonstration that incorporating spatial information can enhance the statistical efficiency of prediction sets. Building on this insight, we propose SACP, a framework that leverages the inherent spatial information in HSIs to improve the performance of conformal prediction. We theoretically demonstrate the validity of SACP by confirming its coverage guarantee.
@@
  • Extensive experiments validate the effectiveness of our proposed method. The emprical results confirm the validity of conformal prediction for HSI classification. Moreover, SACP outperforms standard conformal prediction by reducing the average prediction set size while maintaining satisfactory coverage rates across different scenarios.
@@ II-A Conformal Prediction
$\hat{\pi}_{\theta,y}(\bm{x})=\sigma(f_{\theta}(\bm{x}))_{y}=\frac{e^{f_{\theta,y}(\bm{x})}}{\sum_{i=1}^{\mathit{K}}e^{f_{\theta,i}(\bm{x})}},$
@@
Pipeline
Instead of predicting a single label based on the output probabilities of a classifier, conformal prediction constructs a set-valued mapping $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ that satisfies the following marginal coverage property:
$\mathbb{P}(\mathit{Y}\in\mathcal{C}(\mathit{X}))\geq 1-\alpha,$
@@
Before deployment, conformal prediction starts with a calibration step using a separate calibration set, $\mathcal{D}_{\text{cal}}:=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, drawn independently from the distribution $\mathcal{P}_{\mathcal{XY}}$. Specifically, we compute the non-conformity score $s_{i}=\mathit{S}(\bm{x}_{i},y_{i})$ for each sample in $\mathcal{D}_{\text{cal}}$, which measures how $y_{i}$ “conforms” to the prediction at $\bm{x}_{i}$. The $1-\alpha$ quantile of the scores $\{s_{i}\}_{i=1}^{n}$ is then determined as the threshold $\tau$. Formally, $\tau$ is obtained as follows:
@@
where $\mathcal{C}_{1-\alpha}(\bm{x}_{n+1};\tau)$ includes all labels for which their non-conformity score $\mathit{S}(\bm{x}_{n+1},y)$ does not exceed the threshold $\tau$. As stated in the following theorem, the prediction sets generated by Eqs. (2) and (3) are guaranteed to achieve finite-sample marginal coverage of the true labels.
@@
Notably, the validity is ensured in a distribution-free manner, offering explicit, non-asymptotic guarantees without relying on distributional or model assumptions. Consequently, as long as the exchangeability assumption holds, the generated prediction sets will have a coverage guarantee. This highlights the flexibility of conformal prediction, as it can be applied to any pre-trained classifier when the assumption is satisfied.
@@
Once the base classifier is determined, the effectiveness of the resulting prediction set depends critically on the chosen non-conformity score function $S$. Two popular options are Adaptive Prediction Sets (APS) and Regularized Adaptive Prediction Sets (RAPS). APS calculates the non-conformity scores by accumulating the softmax probabilities, sorted in descending order:
$S_{\text{APS}}(\bm{x},y)\!=\!\sum_{y_{i}\neq y}\hat{\pi}_{\theta,y_{i}}(\bm{x})\cdot\mathds{1}_{\{\hat{\pi}_{\theta,y_{i}}(\bm{x})>\hat{\pi}_{\theta,y}(\bm{x})\}}+u\cdot\hat{\pi}_{\theta,y}(\bm{x}),$
where $u$ is an independent random variable following a uniform distribution on $[0,1]$. However, the softmax probabilities typically exhibit a long-tailed distribution, which facilitates the inclusion of tail classes in the prediction sets. To address this issue, RAPS [33] refines the non-conformity scores by incorporating regularization for the label rankings. Through this regularization, RAPS ensures that labels with higher rankings are associated with larger scores, defined as follows:
$\begin{array}[]{r}S_{\textrm{RAPS}}(\bm{x},y)=S_{\textrm{APS}}(\bm{x},y)+\lambda\cdot(o(y,\hat{\pi}_{\theta}(\bm{x}))-k)^{+},\end{array}$
where $o(y,\hat{\pi}_{\theta}(\bm{x}))$ is the label ranking of $y$, $\lambda$ and $k$ are hyperparameters, and $(z)^{+}$ denotes the positive part of $z$. Yet, the non-conformity score of RAPS still involves unreliable softmax probabilities, leading to suboptimal performance in conformal prediction. To solve it, Sorted Adaptive Prediction Sets (SAPS) [19] reduces the emphasis on probabilities by retaining only the maximum softmax probability and discarding the others. Formally, SAPS is defined as follows:
$S_{\text{SAPS}}(\bm{x},y)=\left\{\begin{array}[]{l}u\cdot\hat{\pi}_{\theta,\max}(\bm{x}),\quad\text{if }o(y,\hat{\pi}_{\theta}(\bm{x}))=1,\\ \hat{\pi}_{\theta,\max}(\bm{x})+(o(y,\hat{\pi}_{\theta}(\bm{x}))-2+u)\cdot\lambda,\text{else},\end{array}\right.$
where $\lambda$ balances the weight of ranking information, $\hat{\pi}_{\theta,\max}(\bm{x})$ denotes the maximum softmax probability of instance $\bm{x}$, and $u$ is a random variable uniformly distributed on $[0,1]$. It can be concluded that these score functions depend exclusively on softmax probabilities and label ranks.
@@
$\text{Coverage}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{(\bm{x},y)\in\mathcal{D}_{\text{test}}}\mathds{1}_{\{y\in\mathcal{C}(\bm{x})\}},$
which should ideally match the pre-defined coverage rate $1-\alpha$. Size measures the average length of prediction sets. When coverage is satisfied, a smaller size indicates higher statistical efficiency and is more desirable. Formally, Size is defined as:
$\text{Size}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{(\bm{x},y)\in\mathcal{D}_{\text{test}}}|\mathcal{C}(\bm{x})|.$
SSCV [33] reflects the conditional coverage rate and is formally defined as:
$\text{SSCV}=100\times\sup_{j}\left|(1-\alpha)-\frac{\left|\{i:y_{i}\in\mathcal{C}(\bm{x}_{i})\text{ and }i\in\mathcal{J}_{j}\}\right|}{|\mathcal{J}_{j}|}\right|,$
@@
II-B HSI Classification
Let $\mathcal{I}\in\mathbb{R}^{H\times W\times U}$ be the HSI to classify, where $H$ and $W$ denote the height and width in pixels, and $U$ denotes the number of spectral bands. Let $M=HW$ be the total number of pixels. Typically, HSI classification aims to predict the label $y\in\mathcal{Y}$ for each pixel in the image, given that a random subset of labels known. The pixels with unknown labels appear randomly at any location in the HSI. When employing a deep classifier, the $(i,j)$-th pixel is generally considered as the center, with information from the small cuboid $B\in\mathbb{R}^{P\times P\times U}$ utilized to extract the joint spectral-spatial feature representation for the central pixel [34] [^1] In prior research, the terms “cuboids”, “cubes”, and “patches” are often used interchangeably to describe local regions of HSIs. For clarity, this work consistently uses the term “cuboids”.. Here, $P$ denotes the patch size of $B$. Correspondingly, the dataset constructed from $\mathcal{I}$ comprises $M$ cuboids, defined as follows:
$\mathcal{D}_{\mathcal{I}}:=\{(B_{i},y_{i})\mid 1\leq i\leq M\},$